4. Find the first order current iſt) in the circuit below it) 5 H 10 A 10 F 3Ω
Answer #1

Given! cit) SH TOA Lof t-D 33 Solution for times it = 0 sec (before switching the capacitor is fully charged and act as an opit for it to А ju SH LOA (1 lof 32 B C Applying KvL in A B C D loop VA - Lolit) - 34t) -Up =0 at VA-VB = Ldlit +3 ict) -O at11 In (it) - ln to -3 S -01 2 ( A ) = In elo = 10A 10A) to te = { from ega [ie lic -0.64] A Ang - 3+ = loe 3 = Loe